# Towards Optimal Distributed Edge Coloring with Small Palettes

### v1

@@ A.1 Replacing the maximal matching by a $2$-edge ruling set
Appendix B Subroutines
B.1 Ruling sets
Theorem B.1 (ruling set algorithm [17, Theorem 8]).
Theorem B.2 (ruling set algorithm [41, Theorem 3]).
Corollary B.3.
Proof B.4.
We start by computing a $\mathcal{O}(\Delta^{2})$-vertex coloring $\varphi$ of $G$ in $\mathcal{O}(\log^{\star}n)$ rounds using Linial’s algorithm [32, 33]. Then, we apply Theorem B.2 to $\varphi$ for $c=\mathcal{O}(\log\Delta)$ to get a $(2,\mathcal{O}(\log\Delta))$-ruling set in time $\mathcal{O}(c\cdot d^{1/c})=\mathcal{O}(\log\Delta\cdot\Delta^{2/\log\Delta})=\mathcal{O}(\log\Delta)$.
Corollary B.5.
@@ B.1 Ruling sets
Proof B.6.
If $\Delta=\mathcal{O}(\log n)$, we can simply use the deterministic algorithm from Corollary B.3 to get an $(2,\mathcal{O}(\log\Delta))$-ruling set (which is also a $(2,\mathcal{O}(\log\log n))$-ruling set) in $\mathcal{O}(\log\Delta+\log^{\ast}n)=\mathcal{O}(\log\log n)$ rounds. Otherwise, for $\Delta=\Omega(\log n)$ we first apply Theorem B.1 to get a $(1,\mathcal{O}(\log\log\Delta))$-ruling set $S_{0}$ with induced degree at most $\mathcal{O}(\log^{5}n)$. Next we compute a $\mathcal{O}(\log^{10}n)$-coloring $\chi$ of $G[S_{0}]$ using Linial’s algorithm [32, 33]. Finally using $\chi$ as the input coloring and setting $c=\mathcal{O}(\log\log n)$, we get a $(2,\mathcal{O}(\log\log n))$ ruling set $S_{1}$ for $S_{0}$ in
@@
Theorem B.7 ($2$-ruling edge set algorithm [31, Theorem 1.3]).
@@
B.2 Hypergraph sinkless orientations
Theorem B.8 (deterministic HSO [8, Theorem 1.4]).
@@ B.2 Hypergraph sinkless orientations
Theorem B.9 (randomized HSO [8, Theorem 1.5]).
@@
B.3 MIS & colorings
Theorem B.10 (deterministic MIS [20, Theorem 1.3]).
@@ B.3 MIS & colorings
Corollary B.11 (randomized MIS [20, Theorem 1.3] & [10]).
@@
Theorem B.12 (deterministic $(3\Delta/2)$-edge coloring [8, Theorem 1.2]).

### v2

@@ A.1 Ruling sets
We start by computing a $\mathcal{O}(\Delta^{2})$-vertex coloring $\varphi$ of $G$ in $\mathcal{O}(\log^{\star}n)$ rounds using Linial’s algorithm [54, 55]. Then, we apply Theorem A.2 to $\varphi$ for $c=\mathcal{O}(\log\Delta)$ to get a $(2,\mathcal{O}(\log\Delta))$-ruling set in time $\mathcal{O}(c\cdot d^{1/c})=\mathcal{O}(\log\Delta\cdot\Delta^{2/\log\Delta})=\mathcal{O}(\log\Delta)$.
Corollary A.5.
@@
Proof A.6.
If $\Delta=\mathcal{O}(\log n)$, we can simply use the deterministic algorithm from Corollary A.3 to get an $(2,\mathcal{O}(\log\Delta))$-ruling set (which is also a $(2,\mathcal{O}(\log\log n))$-ruling set) in $\mathcal{O}(\log\Delta+\log^{\ast}n)=\mathcal{O}(\log\log n)$ rounds. Otherwise, for $\Delta=\Omega(\log n)$ we first apply Theorem A.1 to get a $(1,\mathcal{O}(\log\log\Delta))$-ruling set $S_{0}$ with induced degree at most $\mathcal{O}(\log^{5}n)$. Next we compute a $\mathcal{O}(\log^{10}n)$-coloring $\chi$ of $G[S_{0}]$ using Linial’s algorithm [54, 55]. Finally using $\chi$ as the input coloring and setting $c=\mathcal{O}(\log\log n)$, we get a $(2,\mathcal{O}(\log\log n))$ ruling set $S_{1}$ for $S_{0}$ in
@@
Theorem A.7 ($2$-ruling edge set algorithm [52, Theorem 1.3]).
@@
A.2 Hypergraph sinkless orientation
Theorem A.8 (deterministic HSO [15, Theorem 1.4]).
@@ A.2 Hypergraph sinkless orientation
Theorem A.9 (randomized HSO [15, Theorem 1.5]).
@@
A.3 Maximal Independent Set
Theorem A.10 (deterministic MIS [37, Theorem 1.3]).
@@ A.3 Maximal Independent Set
Corollary A.11 (randomized MIS [37, Theorem 1.3] & [20]).
@@
A.4 Edge coloring
Theorem A.12 (deterministic $(2\Delta-1)$-edge coloring [4, Theorem D.4]).
The $(\mathrm{deg}(e)+1)$-list edge coloring problem can be solved in $\mathcal{O}(\log^{7}C\cdot\log^{5}n+\log^{\ast}n)$ deterministic rounds in the $\mathsf{LOCAL}$ model, where $C$ is the size of the color space.
Theorem A.13 (deterministic $(3\Delta/2)$-edge coloring [15, Theorem 1.2]).